f(x)=6x/x-2, determine the x coordinate of a point (x,y) where 3
Answer by stanbon(75887)

f(x)=6x/x-2, determine the x coordinate of a point (x,y) where 3 ----
Slope of the secant = (18-10)/(3-5) = = -4
----
f'(x) = [(x-2)*6-(6x)(1)]/(x-2)^2
----
Solve f'(x) = -4
----
[(x-2)*6-(6x)(1)]/(x-2)^2 = -4
----
[6x-12-6x] = -4(x-2)^2
---
-12 = -4(x-2)^2
(x-2)^2 = 3
---
x-2 = sqrt(3) or x-2 = -sqrt(3)
x = 2+sqrt(3) or x = 2-sqrt(3)
----
Solution in (3,5) is x= 2+sqrt(3)
================================
Cheers,
Stan H.